OpenStudy (anonymous):

Write the solution set using interval notation. 1/2 (2x+10) > 3/4 (x-1) the answer is (-23, infinity) but how?

OpenStudy (anonymous):

So you want to know how to work out the equation?

OpenStudy (anonymous):

Si

OpenStudy (anonymous):

If so, start by distributing the 1/2 and the 3/4 :)

OpenStudy (anonymous):

I will work it out then

OpenStudy (anonymous):

1/2 (2x+10) > 3/4 (x-1) = x+5> 3/4x-3/4= -3/4x 1/4x +5> -3/4= -5 -5 1/4x= -23/4 = divide by 1/4 or multiply by 4 x=-23

OpenStudy (anonymous):

now, the interval notation you showed suggests that that -23 isn't an included value so the final interval notation will be (-23,infinity)

OpenStudy (anonymous):

C: I hope you understood! ◕ ‿ ◕

OpenStudy (anonymous):

Yes I did thank-you and you are awesome. You make movies and you're a musician like me. O.o

OpenStudy (anonymous):

how did you know? did I put that in my description?

OpenStudy (anonymous):

Yes. What genre of movies do you do?

OpenStudy (anonymous):

Check out Wongfu, that how I roll ◡‿◡

OpenStudy (anonymous):

Umm okay, I film romance. Big time romantic.

OpenStudy (anonymous):

:O oh. Well I cover that too. But, in a single girl kind of perspective

OpenStudy (anonymous):

Big time romantic, but no romance. ha Lets see, The wind blows through Katherine’s smooth hair This makes me stop and stare Her emerging lips painted by an artist’s pen She is a rare rose among many flowers

OpenStudy (anonymous):

Rare rose? I do doubt that, for you see... I have an Identical twin sister 0~0

OpenStudy (anonymous):

and, my hair is very very short. I'll post a new photo

OpenStudy (anonymous):

Yes a rare rose.

OpenStudy (anonymous):

Yays!

OpenStudy (anonymous):

Well that was a fail xD

OpenStudy (anonymous):

You have a tumblr?

OpenStudy (anonymous):

That's a nice pic

OpenStudy (anonymous):

Nope :D Well I do but I don't use it often. Anyways, I have a feeling that I'm much much much much much your senior

OpenStudy (anonymous):

Haha Ursäkta?

OpenStudy (anonymous):

I mean pardon

OpenStudy (anonymous):

I'm betting that I'm older 0_____0

OpenStudy (anonymous):

Haha how old?

OpenStudy (anonymous):

no no no no no, men first

OpenStudy (anonymous):

Ladies first, born a raised to be chivalrous

OpenStudy (anonymous):

So please ladies first

OpenStudy (anonymous):

That's not, at least when a "lady" asks a supposed gent to go first. Also, age is a sketchy thnig when it comes to women :>

OpenStudy (anonymous):

*thing

OpenStudy (anonymous):

Is it between 18-40?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

22-30?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Younger or older?

OpenStudy (anonymous):

you tell me

OpenStudy (anonymous):

your age

OpenStudy (anonymous):

You look young, but you could just look young for your age. Like my grandma doesn't look a day over 30 but she's 79, but that is just the Swede in her.

OpenStudy (anonymous):

And you? *__* 18-25?

OpenStudy (anonymous):

Between 18-70 ha 18-25?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

how old are you now?

OpenStudy (anonymous):

Yeah I'm older. Between 18-20

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Nope ^^

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

Yes indeed

OpenStudy (anonymous):

I am turning 19 in July, so that's not bad. I iz not a coogah. Yay

OpenStudy (anonymous):

Haha that's always a good thing. I think :P

OpenStudy (anonymous):

Agreed.

OpenStudy (anonymous):

So I'm Josh by the way

OpenStudy (anonymous):

Hello Josh. I have a good friend by that name. My full name is Katherine.

OpenStudy (anonymous):

Pleasure to meet you Katherine.

OpenStudy (anonymous):

Same on my part. It does however, feel as if we're playing a game of flattery :)

OpenStudy (anonymous):

Oh does it? c:

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Hmmm I wonder why.

OpenStudy (anonymous):

Because you're a male, and I'm a female. ;} you get it lass

OpenStudy (anonymous):

Yes :3

OpenStudy (anonymous):

So do you sing?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

:D I used to, what do you sound like?

OpenStudy (anonymous):

a girl ;]

OpenStudy (anonymous):

I sound like Brooke Fraser

OpenStudy (anonymous):

No! Really?! I did not know haha Ohh

OpenStudy (anonymous):

Yeeeup. But I actually can sing though xD

OpenStudy (anonymous):

Do you have like a clip you recorded?

OpenStudy (anonymous):

Nope. And I'm doing homework right now.

OpenStudy (anonymous):

it's late too xD 11pm almost

OpenStudy (anonymous):

Oh ok, want me to let you go? Its like 1am here

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I'll be back on in the morining :) around 9am

OpenStudy (anonymous):

Talk to you then ^^

OpenStudy (anonymous):

Ok, want my email cause I would like you to look over my poems

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

callisjoshua1@gmail.com

OpenStudy (anonymous):

hey hey, :) I'm late. I'll message you later tonight. I'm swamped with math :O

OpenStudy (anonymous):

Hey hey, Okay sounds good

OpenStudy (anonymous):

Hiya :] I've been very busy with classes, sorry I haven't been on!

OpenStudy (anonymous):

Its fine c: